OpenStudy (anonymous):

Which of the following describes the non-rigid transformation in the function shown below? y-1=-(3x+1)^2 A. The graph is shifted 3 units down. B. The graph is stretched vertically by a factor of 3. C. The graph is reflected across the x-axis. D. The graph is stretched horizontally to 1/3 the original width.

OpenStudy (anonymous):

The first question you need to ask yourself is what is the basic function? As in the type of function you're working with.

OpenStudy (anonymous):

would it be point-slope?

OpenStudy (anonymous):

try to get it to slope intercept form y=mx+b

OpenStudy (anonymous):

@ali1029 you there O.O

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

y-1=-(3x+1)^2 y-1=-3x-1^2 y-1=-3x1 y=-3x+2

OpenStudy (anonymous):

No I was more referring to the fact that we are dealing with a quadratic function. As in y(x)=ax^2.

OpenStudy (anonymous):

ohhh you should listed to him he seems better at this than me

OpenStudy (anonymous):

Solving for y yields:\[y=-(3x+1)^{2}+1\]Factoring out the 3:\[y=-(3(x+\frac{ 1 }{ 3 }))^{2}+1\] From this we can see that the base function y(x)=x^2 has been: 1)stretched vertically by a factor of 3: y(3x)=(3x)^2 2)shifted left 1/3 units: y(x+1/3)=(3(x+1/3))^2 3)flipped across the y axis: -y(x)=-(3(x+1/3))^2 4)shifted down 1 unit y(x)-1=-(3(x+1/3))^2 which is what we started out with.

OpenStudy (anonymous):

I hope that isn't too confusing. I'm trying to show how this is really just taking steps and building up the function based on the changes you want to make to the base function.

OpenStudy (anonymous):

im still trying to understand this, sorry

OpenStudy (anonymous):

would the answer be B? because it is going up higher and vertically?

OpenStudy (anonymous):

Give me your number and, I'll text you the answer.

OpenStudy (anonymous):

Creepy

OpenStudy (anonymous):

Just one second. I'm making sure I didn't switch the terminology for the stretching part.

OpenStudy (anonymous):

Get out @naevusx . That is really creepy.

OpenStudy (anonymous):

& okay

OpenStudy (anonymous):

Actually the step "1)stretched vertically by a factor of 3: y(3x)=(3x)^2" should be shrunk horizontally by a factor of 1/3.

OpenStudy (anonymous):

It's because you're causing the function to reach y values 3 times as fast so it kind of compresses the x axis if that makes sense.

OpenStudy (anonymous):

oh okay, i see what you're saying

OpenStudy (anonymous):

 kind of like this